Calculate the number of grams of aluminum sulfate produced if three point 85 grams of aluminum is added to six point 85 grams of sulfuric acid.
1
Expert's answer
2021-10-17T11:20:51-0400

2Al(s)+3H2SO4(aq)→Al2(SO4)3(aq)+3H2(g)

Molar mass of Al= 26.981539

85/26.981539= 3.15mol

Molar mass of H2SO4= 98.079

85/98.079= 0.87mol

Mole ratio of Al:Al2(SO4)3= 2:1

Mole ratio of H2SO4:Al2(SO4)3= 3:1

= 1/2+1/3= 5/6

Molar mass of Al2(SO4)3= 342.15 g/mol

342.16×5/6

= 285.13g